# One Wave to Explain Them All: A Unifying Perspective on Post-hoc Explainability

### v1

@@ 4.2 XAI Properties of the Wavelet Attribution Method
Figure 4: Qualitative illustration of WAM for audio via a Noise experiment. We add 0 dB white noise on he audio of the target class (’Crow’) to form the input to the classifier. Interpretation audio reconstructed with important wavelet coefficients virtually eliminate noise, and also clearly emphasize parts of the target class audio (indicated with green box).
Explanations on voxels.
We retrieve on voxels the same decomposition as for images or audios. Figure 5 highlights the significance of the edges at larger scales, whereas at smaller scales, the importance becomes increasingly concentrated at the center of the digit. To the best of our knowledge, WAM is the first method to show such decomposition on shapes.
Figure 5: Decomposition of the different important scales on a voxel with the GradWCAM.
@@ Conclusion.
We have introduced a novel approach for feature attribution by computing explanations in the wavelet domain rather than the input domain, providing a framework applicable to audio, images, and shapes. This method shifts away from traditional pixel-based decompositions used in saliency mapping, offering more precise insights into model decisions by leveraging the wavelet domain’s ability to preserve inter-scale dependencies. This ensures that critical aspects like frequency and spatial structures are maintained, resulting in richer explanations compared to traditional feature attribution methods.
Our method, WAM, shows a strong ability to highlight essential audio components in noisy samples, isolate necessary shape and texture features for accurate predictions, and offer richer explanations for shape classification. Quantitatively, it achieves state-of-the-art results across both audio and image benchmarks.
@@
Despite its advantages, the current method does not extend to 3D point cloud data, and for audio, the greedy extraction of important coefficients is unsuitable for generating listenable explanations. Future work could explore alternative wavelet decompositions, such as continuous or complex wavelets for audio explanations and graph wavelet transforms to handle unstructured point clouds. Additionally, our method could be applied to videos mathematically similar to the voxel data used in this work. We hope this approach will inspire further research into the properties of explanation domains, the wavelet domain being one such domain.
Appendix A Implementational details
A.1 Effect of smoothing and integration on the explanations
Figure 6 illustrates the improvement of the quality of the explanations by smoothing (third row) or integrating (fourth row) the gradients, compared to their raw values (second row). Smoothing follows Equation 2 and integration follows Equation 3.
@@ A.1 Effect of smoothing and integration on the explanations
Figure 6: Effect of smoothing (3rd row) and integration (4th row) compared to the raw gradients (2nd) raw when computing the WAM of the images (1st row). The explanations are depicted in the wavelet domain.
We can see that both methods display complementary properties regarding the explanation. WAMSG enables to visualize highlights the important locations within scales, while WAMIG emphasizes on see the relative importance of each scale.
A.2 Benchmark construction
All benchmarks reported in this work were carried out on a server running Ubuntu 20.04.6 and on a single NVIDIA TITAN Xp GPU with 12 GB of VRAM with CUDA 12.5. The data to replicate the experiments can be downloaded in the repository accessible at this URL https://doi.org/10.5281/zenodo.13873810, and the source code is accessible from the Git repository.
@@ A.2 Benchmark construction
  • ResNet: we consider the resnet18 variant,
  • EfficientNet: we consider the tf_efficientnet_b0.ns_jft_in1k variant,
  • ConvNext: we consider the convnext_small.fb_in22k_ft_in1k_384 variant,
  • DeiT: we consider the deit_tiny_patch16_224.fb_in1k
@@
Figure 7: Pseudo-code for adding Gaussian noise to audio
Appendix B Complements on the quantitative evaluation
B.1 Insertion and deletion
Insertion and deletion are two evaluation metrics proposed by Petsiuk et al. (2018). These metrics are ”area-under-curve” (AUC) metrics, which report the change of in the predicted probability for the image class when inserting (resp. removing) meaningful information highlighted by the attribution method. Petsiuk et al. (2018) initially defined this metric for images, where the important features correspond to pixels. We expand this metric to wavelet coefficients, thus enabling a computation of the Insertion and the Deletion for any modality.
@@ B.2.1 Definitions
The Faithfulness on Spectra (FF, Parekh et al., 2022) measures how important is the generated interpretation for a classifier. The metric is calculated by measuring the drop in class-specific logit value $\bm{f}(\bm{x})_{c}$, when the masked out portion of the interpretation mask $\bm{m}_{\bm{\gamma}}$ is input to the classifier. This amounts to calculating,
@@
B.2.2 Results
Images.
Table 3 reports the results using the $\mu$-Fidelity (Bhatt et al., 2021) compared to other methods. We can see that the performance measured by the $\mu$-Fidelity is more in line between our method and the existing approaches.
Table 3: $\mu$-Fidelity (Bhatt et al., 2021) score obtained on 1,000 images from the validation set of ImageNet and for different model architectures. Higher is better. Bolded results are the best and underlined values are the second best.
Method ResNet-18 ConvNext EfficientNet DeiT Mean
Saliency 0.154 0.186 0.180 0.195 0.179
Integrated Gradients 0.228 0.223 0.219 textbf0.241 0.228

### v2

@@ 4.2 Exploring how WAM renews feature attribution
Figure 5: Qualitative illustration of WAM for audio in a noise experiment. We add 0 dB white noise to the audio of the target class (“Crow”) as input to the classifier. Audio reconstructed using important wavelet coefficients effectively removes noise and highlights key parts of the target class (marked with a green box). WAM identifies these relevant parts in a fully post-hoc manner.
@@ Conclusion.
We bring a novel perspective on feature attribution by computing explanations in the wavelet domain rather than the pixel domain, providing a framework applicable to audio, images, and volumes. This method shifts away from traditional pixel-based decompositions used in saliency mapping, offering more precise insights into model decisions by leveraging the wavelet domain’s ability to preserve inter-scale dependencies. This ensures that critical aspects like frequency and spatial structures are maintained, resulting in richer explanations compared to traditional feature attribution methods.
Our method, WAM, shows a strong ability to highlight essential audio components in noisy samples, isolate necessary volumes and texture features for accurate predictions, and offer richer explanations for shape classification. Quantitatively, it achieves state-of-the-art results across audio, image, and volume benchmarks.
@@
Despite its advantages, the current method does not extend to point cloud data. For audio, the greedy extraction of important coefficients is unsuitable for generating listenable explanations.
@@
Future work could explore alternative wavelet decompositions, such as continuous or complex wavelets for audio explanations and graph wavelet transforms to handle unstructured point clouds. Additionally, our method could be applied to videos, mathematically similar to the 3D data used in this work.
The unification of the feature attribution domain finds a natural application in explaining multi-modality models (e.g., Vision-Langage Models). Formally, the wavelet decomposition cannot be applied to text data, thus limiting the applicability of our method in natural language processing applications. Solutions to overcome this difficulty may be found by building on existing methods in this field, which have been surveyed by Lyu et al. (2024).
More broadly, we hope that this work will open the discussion on the properties and expressiveness of the domain in which feature attribution is carried out. In particular, the theoretical properties of expanding feature attribution beyond the pixel space could be further discussed.
Impact Statement
This work addresses the growing need for grounded, interpretable AI systems, especially in light of increasing regulatory demands for transparency in machine learning. By providing a unified approach to explainability across diverse modalities (audio, image, and volumes) our method ensures that model explanations remain faithful to the inherent structure of the data. In parallel, as multimodal models become more prevalent, our work offers a solution that can be applied to any input type, facilitating more consistent and robust explanations.
Appendix A Implementational details and consistency checks
A.1 Effect of smoothing and integration on the explanations.
Figure 6 illustrates the improvement of the quality of the explanations by smoothing (third row) or integrating (fourth row) the gradients, compared to their raw values (second row). Smoothing follows Equation 7 and integration follows Equation 8. We can see that both methods display complementary properties regarding the explanation. WAMSG enables to visualize highlights the important locations within scales, while WAMIG emphasizes on the relative importance of each scale. This suggests that the variants of WAM can be used for different applications. If one is interested in studying the inter-scale dependencies, the reliance on different frequency ranges, or more broadly, the robustness of the model, then WAMIG should be preferred. On the other hand, if one wants to focus on the important details at each scale, for instance, to highlight the important patterns for the prediction of a given class, then WAMSG appears as more relevant.
@@ Appendix A Implementational details and consistency checks
Figure 6: Effect of averaging or smoothing (3rd row) and integration (4th row) compared to the raw gradients (2nd) when computing the WAM of the images (1st row). The explanations are depicted in the wavelet domain. ”Smoothing” corresponds to the application of SmoothGrad (Smilkov et al., 2017) and ”Integrating” to the application of Integrated Gradients (Sundararajan et al., 2017).
A.2 Effect of the mother wavelet
Definition.
The mother wavelet is the fundamental function $\bm{\psi}\in L^{2}(\mathbb{R}^{n})$ used to generate a family of wavelets through dilation and translation. This construction enables multiresolution analysis. From a mother wavelet $\bm{\psi}(\bm{x})$, the family of function is defined as
$\psi_{\bm{\lambda},\bm{b}}(\bm{x})=\frac{1}{\bm{\lambda}^{n/2}}\left(\frac{\bm{x}-\bm{b}}{\bm{\lambda}}\right),$ (10)
where $\bm{\lambda}\in\mathbb{R}^{n}\setminus\{\mathbf{0}\}$ and $\bm{b}\in\mathbb{R}^{n}$ are respectively the scale and translation parameters. The choice of the mother wavelet influences which features are captured, the quality of signal reconstruction, and the effectiveness of tasks like compression, denoising, and feature extraction.
Interpretability.
WAM formally indicates that the model is sensitive to a given region in the space-scale domain. Since the reconstructed image is the same, no matter the chosen basis for decomposition, the interpretation in terms of structural components (i.e., what those coefficients correspond to) depends on the choice of the mother wavelet. The choice of the mother wavelet is therefore grounded in the application of interest, and the information that is invariant to the choice of the mother wavelet is the importance of of a given scale at a given location.
Among the popular mother wavelet, the Haar wavelet (Haar, 1910) enables us to interpret the sensitivity of the model in terms of sharp changes, edges, or discontinuities on the image. It can be used for instance to detect grid-like structures. Daubechies wavelets (Daubechies, 1988) are smoother than Haar wavelets and highlight smooth patterns or fine textural details. Finally, the Biorthogonal (bior) wavelet (Cohen et al., 1992) underlines curves, contours, and symmetric structure, especially on images.
Figure 7 illustrates the different structural features emphasized depending on the choice of the mother wavelet. In this example, we can see that the Haar wavelet decomposes the image in terms of sudden transitions, highlighting the edges of the input image. In contrast, using the Daubechies wavelet reveals more of the image textures. In Section D.4 we illustrate the usefulness of WAM in a practical setting, where the chosen mother wavelet was the Haar wavelet, since the object of interest was to understand the model’s behavior with respect to the detection of sharp objects on images.
Figure 7: Impact of the mother wavelet. Different mother wavelets emphasize different features of the image.
Quantitative checks.
Table 2 empirically checks that the explanations generated by WAM are invariant to the choice of the mother wavelet. For images and volumes, the insertion and deletion scores remain unchanged. Results are reported up to the 3rd decimal, but results remain the same up to the 8th decimal. For audio, the numerical results are close but not identical. This slight variation arises because the reconstruction of the input signal depends on the choice of the mother wavelet. Consequently, different wavelets can lead to minor differences in the results. However, when the reconstructions align, the resulting explanations are identical.
Table 2: Impact of the mother wavelet. Changing the mother wavelet does not affect the WAMIG Insertion and Deletion scores for images and volumes, but it causes slight variations for audio.
Audio Images Volumes
Model ResNet ResNet ResNet
Dataset ESC-50 ImageNet AdrenalMNIST
Insertion Deletion Insertion Deletion Insertion Deletion
Bior 0.442 0.256 0.422 0.078 0.404 0.631
Daubechies 0.446 0.266 0.422 0.078 0.404 0.631
Haar (baseline) 0.438 0.263 0.422 0.078 0.404 0.631
A.3 Randomization checks
Motivation.
The sanity checks introduced by Adebayo et al. (2018) aim at assessing whether an explanation depends on the model’s parameters and the input labels. These tests assess the faithfulness of an explanation beyond visual evaluation. The randomization test evaluates whether an explanation depends on the model’s parameters. Parameters have a strong effect on a model’s performance. Therefore, for a saliency method to be useful for debugging or analyzing a model, it should be sensitive to its parameters. Adebayo et al. (2018) proposed different methods to randomize the model parameters. One particularly interesting implementation is the “cascading” randomization, in which the weights are randomized from the top to the bottom layers.
Method.
We compute WAM for 1,000 ImageNet validation samples for a set of increasingly randomized models. A randomized layer is a layer that we reset at its initial value. We consider a ResNet-18 and randomize its layers from the shallowest conv1 to the deepest fc. We then compute the rank correlation (or Pearson correlation coefficient, Pearson, 1896) between the WAM of the original, fully-trained model (labeled orig) and the randomized models (labeled by the name of the layer until which they are randomized).
Results.
Figure 8 presents the results. The dotted line represents the average rank correlation across the 1,000 images, and the intervals represent the 95% confidence intervals. We can see that the correlation between WAM s significantly decreases as the randomization increases, thereby showing that WAM is sensitive to the model’s parameters. The lower decrease that we observe for WAMIG compared to WAMSG comes from the fact that WAMIG reflects more the inter-scales distribution of the importance than WAMSG does. As pointed out by Rahaman et al. (2019) and Yin et al. (2019), even random models exhibit a spectral bias, i.e., a natural tendency to favor lower frequencies over higher ones, which translates here into the fact of naturally putting more importance on coarser scales rather than finer scales, no matter the depth of the randomization. Figure 9 qualitatively illustrates how the explanations evolve as the randomization depth increases.
Appendix B Theoretical properties
Sundararajan et al. (2017) highlighted two axioms that feature attribution methods should satisfy, namely sensitivity and implementation invariance. We first recall the definitions of these axioms.
Axiom B.1 (Sensitivity (a)).
If two inputs $\bm{x}$ and $\bm{x}^{\prime}$ differ in one feature and have different predictions, then the differing feature must receive a non-zero attribution.
Axiom B.2 (Implementation Invariance).
Two functionally equivalent networks (i.e., computing the same function) must yield the same attributions.
Since WAMIG builds on integrated gradients, we may wonder whether it inherits the properties of the original integrated gradients. In the following, let $\bm{f}_{c}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ be the class logit function in input space $\mathcal{W}$ denotes the wavelet transform operator and $\mathcal{W}^{-1}$ its inverse. We define $\bm{z}=\mathcal{W}(\bm{x})$ and let $\bm{z}_{0}$ be a fixed baseline in the wavelet space. We also define the composed function $f=\bm{f}_{c}\circ\mathcal{W}^{-1}$, such that $f(\bm{z})=\bm{f}_{c}(\bm{x})$.
B.1 WAMIG satisfies completeness
We can show that under some conditions on the choice of the mother wavelet $\bm{\psi}$, WAMIG satisfies completeness, and thus sensitivity. Following Sundararajan et al. (2017), this implies that WAMIG  satisfies sensitivity (a).
Definition B.1 (Completeness).
An attribution method $\bm{\gamma}$ satisfies completeness if
$\sum_{i=1}^{n}\bm{\gamma}(\bm{x})_{i}=F(\bm{x})-F(\bm{x}_{0}).$ (11)
for any given $F:\mathbb{R}^{n}\to\mathbb{R}$ differentiable almost everywhere.
Theorem B.1.
If $\mathcal{W}^{-1}$ is differentiable almost everywhere, then WAMIG satisfies completeness:
[…]
Proof.
Proposition 1 in Sundararajan et al. (2017) states that for some $f$ differentiable almost everywhere,
$\sum_{i=1}^{n}IG_{i}(\bm{x})=f(\bm{x})-f(\bm{x}_{0})$
since by definition, WAMIG is IG applied to $f(\bm{z})$, we have
[…]
Now taking $f=\bm{f}_{c}\circ\mathcal{W}^{-1}$, Equation $\star$ can be re-written as
[…]
$=\bm{f}_{c}(\mathcal{W}^{-1}(\bm{z}))-\bm{f}_{c}(\mathcal{W}^{-1}(\bm{z}_{0}))$
$=\bm{f}_{c}(\bm{x})-\bm{f}_{c}(\bm{x}_{0})$
and $\bm{f}_{c}\circ\mathcal{W}^{-1}$ remains differentiable (almost everywhere) provided $\mathcal{W}^{-1}$ is differentiable. This can be enforced by choosing a mother wavelet $\bm{\psi}$ that is smooth (e.g., with Daubechies but not with Haar wavelets). Thus, provided $\bm{\psi}$ is smooth, WAMIG  satisfies completeness. ∎
B.2 Additional theoretical properties
It is also possible to show that WAMIG satisfies implementation invariance and linearity. As with completeness, it is enough to show that the corresponding composed function $f$ “respects” implementation invariance or linearity of $\bm{f}_{c}$. And since $IG$ satisfies both these properties, WAMIG will satisfy as well as it is IG applied to the corresponding $f$ functions.
For implementation invariance, if $\bm{f}_{c}^{(1)},\bm{f}_{c}^{(2)}$ have different architectures but exactly same output for any input $x$, the corresponding composed functions $f_{(1)},f_{(2)}$ also have the same output for any input $z$ and thus WAMIG outputs are exactly the same for any given input $x$ for $\bm{f}_{c}^{(1)},\bm{f}_{c}^{(2)}$.
For linearity, one needs to show that WAMIG output for a linear combination of any two functions $\bm{f}_{c}^{(3)}=a\bm{f}_{c}^{(1)}+b\bm{f}_{c}^{(2)}$ is linear combination of WAMIG output for each of them. Let $f_{(i)}$ be the associated composed function of $\bm{f}_{c}^{(i)}$. Then $f_{(3)}=af_{(1)}+bf_{(2)}$ and because of linearity of $IG$ applied on $f_{(i)}$’s, WAMIG satisfies linearity as well. Thus, it’s easy to see that WAMIG also satisfies implementation invariance and linearity axioms. More generally, we believe that WAMIG is a path method, and thus inherits the properties of the path methods proved by Friedman (2004). However, the focus of the paper is on the proposal and practical usage of wavelet attribution methods. We keep any formal study of theoretical properties of WAMIG or other variants of WAM as future work.
B.3 Feature attribution beyond the input space
More generally, our approach expands attribution methods on transformations of the input in an alternative domain. It seems that the properties satisfied by traditional attribution method generalize to the case where attribution is done on transformations of the input data. This work thus opens a promising direction in the systematic study of generalized attribution spaces, or transformed domains where gradients can be computed more expressively while preserving fidelity to the original input.
Appendix C Complements on the quantitative evaluation
C.1 Benchmark construction
@@
  • ResNet (He et al., 2016): unless specified otherwise, we consider the resnet18 variant,
  • EfficientNet (Tan & Le, 2019): we consider the tf_efficientnet_b0.ns_jft_in1k variant,
  • ConvNext (Liu et al., 2022): we consider the convnext_small.fb_in22k_ft_in1k_384 variant,
  • DeiT (Touvron et al., 2021): we consider the deit_tiny_patch16_224.fb_in1k
@@
Figure 10: Pseudo-code for adding Gaussian noise to audio.
Volumes.
For volumes, we consider the LATEC benchmark (Klein et al., 2024) and evaluate our method on two datasets of MedMNIST (Yang et al., 2023): AdrenalMNIST3D and VesselMNIST. We carry out our evaluation over the complete test set. We included WAM  as an additional feature attribution method for fair comparisons with existing methods.
C.2 Definition of the evaluation metrics
Insertion and Deletion.
Insertion and deletion are two evaluation metrics proposed by Petsiuk et al. (2018). These metrics are ”area-under-curve” (AUC) metrics, which report the change in the predicted probability for the image class when inserting (resp. removing) meaningful information highlighted by the attribution method. Petsiuk et al. (2018) initially defined this metric for images, where the important features correspond to pixels. We expand this metric to wavelet coefficients, thus enabling computation of the Insertion and the Deletion for any modality.
@@ C.2 Definition of the evaluation metrics
The Faithfulness on Spectra (FF, Parekh et al., 2022) measures how important is the generated interpretation for a classifier. The metric is calculated by measuring the drop in class-specific logit value $\bm{f}(\bm{x})_{c}$ when the masked out portion of the interpretation mask $\bm{m}_{\bm{\gamma}}$ is input to the classifier. This amounts to calculating,
@@
C.3 Additional evaluation results
This section provides additional quantitative results. For audio, we evaluated WAM using additional metrics and on a noisy version of the dataset ESC-50 and SONY-UST, as done by Parekh et al. (2022). We also evaluated WAM by computing explanations from the mel-spectrogram and the wavelet coefficients of the input audio. For images, we evaluated our method using different model topologies: ResNet, EfficientNet, Data Efficient Transformer, and ConvNext. We also included additional metrics and additional methods. For volumes, we evaluated the performance of our method on MNIST3D using a ResNet3D backbone and we included two additional datasets, AdrenalMNIST and VesselMNIST, and three model topologies: ResNet3D, EfficientNet3D, and 3D Former. We also explored the effect of the number of decomposition levels on performance. Across all these variants, WAM demonstrated solid performance compared to the baselines, showing its robustness and versatility across use cases and datasets.
Audio
Table 3 presents the full evaluation results for all audio metrics. We report the results for the clean and noisy dataset, as done by Parekh et al. (2022). For WAM, we consider two variants: when the explanations are computed on the mel-spectrogram of the input audio and when it is computed from the wavelet coefficients. We can see that when the explanations is computed from the wavelet coefficients, accuracy is higher. We report the results on two datasets, ESC-50 (Piczak, 2015) and SONY-USC (Cartwright et al., 2019)
Table 3: Evaluation scores of WAM and comparison with baselines on 400 audio samples from ESC-50 (fold 1) and on the validation set of the SONY-UST dataset. The column ”-” indicates that the samples are unaltered. The column ”+WN” indicates that the samples have 0 dB Gaussian white noise. Bolded results are the best and underlined values are the second best.
Metric Faithfulness ($\uparrow$) Insertion ($\uparrow$) Deletion($\downarrow$) FF ($\uparrow$) Fid-In ($\uparrow$)
- +WN - +WN - +WN - +WN - +WN
ESC-50 IntegratedGradients 0.264 0.310 0.267 0.312 0.047 0.045 0.207 0.207 0.220 0.225
GradCAM 0.072 0.073 0.274 0.274 0.201 0.201 0.137 0.135 0.517 0.542
Saliency 0.066 0.065 0.220 0.221 0.154 0.156 0.166 0.168 0.253 0.245
WAMSG (ours, wavelet) 0.197 0.205 0.449 0.452 0.252 0.246 0.132 0.130 0.718 0.690
WAMIG (ours, wavelet) 0.176 0.182 0.436 0.442 0.260 0.261 0.118 0.124 0.652 0.657
WAMSG (ours, melspec) 0.009 0.007 0.169 0.166 0.159 0.161 0.152 0.149 0.117 0.122
WAMIG (ours, melspec) 0.000 0.004 0.168 0.171 0.168 0.167 0.149 0.149 0.105 0.128
SONY-UST IntegratedGradients 0.151 0.151 0.732 0.732 0.581 0.582 -0.122 -0.121 0.000 0.002
GradCAM 0.027 0.026 0.600 0.598 0.573 0.571 -0.004 -0.004 0.800 0.774
Saliency 0.000 0.003 0.706 0.708 0.706 0.705 -0.224 -0.225 0.000 0.002
SmoothGrad 0.115 0.116 0.682 0.682 0.567 0.566 -0.098 -0.098 0.000 0.002
WAMSG (ours, wavelet) -0.076 -0.076 0.598 0.600 0.674 0.675 0.029 0.026 0.882 0.878
WAMIG (ours, wavelet) -0.075 -0.080 0.591 0.591 0.666 0.671 0.037 0.036 0.866 0.840
WAMSG (ours, melspec) 0.003 0.010 0.677 0.679 0.674 0.669 -0.188 -0.183 0.000 0.002
WAMIG (ours, melspec) 0.028 0.023 0.728 0.723 0.700 0.700 -0.202 -0.200 0.000 0.000
Images
For images, we report the evaluation results across different model topologies and with additional methods. On a ResNet, we also compare our method with more recent feature attribution methods. All explanations are computed on 1,000 samples from the validation set of ImageNet.
Faithfulness and $\mu$-Fidelity.
Table 4 reports the evaluation results using different metrics across a wide range of topologies. We also include additional methods such as GradCAM ++ (Selvaraju et al., 2017) and Guided Brackpropagation (Springenberg et al., 2014). We can see that in terms of Faithfulness, our method outperforms alternative approaches over all model topologies. On the other hand, the performance measured by the $\mu$-Fidelity aligns WAM with existing approaches. We can see that the good results reported in Table 4 are mostly driven by our method’s very good results in Insertion. We can see that WAM systematically outperforms the other methods in both Insertion and Deletion.
Table 4: Evaluation results of WAM for images. We report the Faithfulness (Muzellec et al., 2024), $\mu$-Fidelity (Bhatt et al., 2021) and Insertion and Deletion (Petsiuk et al., 2018) scores obtained on 1,000 images from the validation set of ImageNet and for different model architectures. Bolded results are the best and underlined values are the second best.
Model ResNet ConvNext EfficientNet DeiT Mean
Faithfulness ($\uparrow$) Saliency 0.025 0.032 0.008 0.038 0.025
Integrated Gradients 0.000 0.001 0.000 0.003 0.001
GradCAM 0.134 0.072 0.061 0.162 0.107
GradCAM++ 0.184 0.055 0.050 0.044 0.083
SmoothGrad 0.023 0.000 0.010 0.004 0.009
Guided-Backpropagation 0.001 0.001 0.001 0.000 0.000
WAMSG (ours) 0.438 0.334 0.350 0.423 0.386
WAMIG (ours) 0.344 0.359 0.370 0.420 0.373
$\mu$-Fidelity ($\uparrow$) Saliency 0.154 0.186 0.180 0.195 0.179
Integrated Gradients 0.228 0.223 0.219 0.241 0.228